OpenStudy (anonymous):

Suppose that w and t vary inversely and that t =1/5 when w = 4. Write a function that models the inverse variation, and find t when w = 9. Would it be t=1/5w;4/45 or 4/5w;4/45?

hartnn (hartnn):

inverse variation ---> wt = constant let wt=k put w=4, t= 1/5 and find k. ?

OpenStudy (anonymous):

4/5? :/

hartnn (hartnn):

yes, wt = 4* (1/5) = 4/5 = k so your function is wt = 4/5 or t= 4/(5w) now to get t, when w=9, just put w=9 in t=4/(5w)

OpenStudy (anonymous):

ooooh, so it would be 4/45?

hartnn (hartnn):

yes, absolutely :)

OpenStudy (anonymous):

your awesome! :D thanks again! xx

hartnn (hartnn):

your awesome too ! :D welcome again ^_^